OpenStudy (anonymous):

9x to the 4th power minus 37x squared plus 4=0 please help

OpenStudy (anonymous):

\[9x^4-37x^2+4=0\]?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

ok trick is to put \[z=x^2\] and rewrite as \[9z^2-37z+4=0\] which by some miracle factors as \[(9z-1)(z-4)=0\] so \[z=\frac{1}{9}\] or \[z=4\]

OpenStudy (anonymous):

then go back and put \[z^2=\frac{1}{9}\] so \[x=\pm\frac{1}{3}\] or \[x^2=4\] so \[x=\pm 2\] and those are your 4 answers.

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

Amazing thank you very much !